OpenStudy (anonymous):

An I.V. solution contains 9.0 mg of sodium chloride per mL. The I.V. solution is being administered at a rate of 1.5 mL per minute. How many grams of sodium chloride would a patient receive in 4.5 hours

OpenStudy (anonymous):

9.0mg * 1g/1000mg * 1.5mL/minute * 60minutes/hour * 4.5hours should be what you/re looking for.

OpenStudy (xishem):

\[\frac{9.0mg(NaCl)}{1mL(NaCl)}*\frac{1.5mL(NaCl)}{1\min}*\frac{60\min}{1hr}*\frac{1g(NaCl)}{1000mg(NaCl)}*\frac{4.5hr}{1}=3.6g(NaCl)\]

OpenStudy (anonymous):

why can you go 1ml/9.0mg? And where did you get the 1 underneath 4.5gr?

OpenStudy (anonymous):

cant*

OpenStudy (xishem):

Because there are 9.0mg of NaCl PER mL of solution. The one means nothing. It's just to make it more readable.

OpenStudy (xishem):

Sorry. To answer your first question, you can use 1mL/9.0mg. I just chose the reciprocal. You will get the same answer either way if you follow the units.